Let fx) = 5x- 5 and g(x) = -2x+9. Find: %D %3D (f-3)(3).
Answer #1

f(n)=5n-5 g(n)- - 2a49 (f-9)(3) = f(3) - 913) f(3) = 5(3)-5 - 15-5 = 10 913) = -2(3) +9 = - 679 = 3 (49)(3) = 10-3 7/{847X1X